In last night’s basketball game, Sally scored 9 less than twice what Lucy scored. The sum of their scores is 27. How many points did Sally and Lucy make individually?
1
Expert's answer
2012-11-29T08:07:04-0500
the equations are
S+L=27
2*L - S = 9
where S is Sally score and L is Lucy score. Hence
S = 2L-9
L + 2L - 9 = 27
3L=36
L=12, S=27-12 = 15
Lucy scored 12, Sally scored 15